Set up a vector n with elements 1, 2, 3, 4, 5. Use MATLAB array opera- tions on it to set up the following four vectors, each with five elements: (a) 2, 4, 6, 8, 10 b)
Answer #1

a = 2:2:10
b = (1/2):(1/2):(5/2)
c = 1./(1:5)
d = 1./((1:5).^2)

\;\;